A 4.5-kg block moving at 2.0 m/s west on a frictionless surface collides totally inelastically with a second 1.0-kg block traveling east at 2.0m/s.

a) Determine the final velocity of the blocks.

b)Determine the kinetic energy of the first block before the collision.

c)Determine the kinetic energy of the second block before the collision.

d)Determine the kinetic energy of the first block after the collision.

e)Determine the kinetic energy of the second block after the collision.
Answer #1

Mass of the first block is m_1=4.5\ kg

Mass of the second block is m_2=1.0\ kg

Initial velocity of the first block v_1=2.0\ m/s

Initial velocity of the second block v_2=-2.0\ m/s

We have taken the velocity in the west direction to be positive and the velocity in the east direction to be negative.

(a)

Initial momentum of the system of two blocks is

Pi= m1v1 + m2uz

P_i=(4.5\ kg)(2.0\ m/s)+(1.0\ kg)(-2.0\ m/s)

P_i=9.0\ kgm/s-2.0\ kgm/s

P_i=7.0\ kgm/s

In totally inelastic collision, the two blocks stick together after the collision, and move together. with the same velocity. Let v be the final velocity of the two blocks. The final momentum of the blocks is

P_f=m_1v+m_2v

P_f=(m_1+m_2)v

P_f=(4.5\ kg+1.0\ kg)v

P_f=(5.5\ kg)v

Applying conservation of momentum

P_f=P_i

(5.5\ kg)v=7.0\ kgm/s

v=\frac{7.0\ kgm/s}{5.5\ kg}

\boxed{v=1.27\ kgm/s}

(b)

Kinetic energy of the first block before the collision is

K_{1i}=\frac12m_1v_{1i}^2

K_{1i}=\frac12\times(4.5\ kg)\times(2.0\ m/s)^2

\boxed{K_{1i}=9.0\ J}

(c)

Kinetic energy of the second block before the collision is

K_{2i}=\frac12m_2v_{2i}^2

K_{2i}=\frac12\times(1.0\ kg)\times(-2.0\ m/s)^2

\boxed{K_{2i}=2.0\ J}

(d)

Kinetic energy of the first block after the collision is

K_{1f}=\frac12m_1v^2

K_{1f}=\frac12\times(4.5\ kg)\times(1.27\ m/s)^2

\boxed{K_{1f}=3.63\ J}

(e)

Kinetic energy of the second block after the collision is

K_{2f}=\frac12m_2v^2

K_{2f}=\frac12\times(1.0\ kg)\times(1.27\ m/s)^2

\boxed{K_{2f}=0.806\ J}